OpenStudy (anonymous):

ABCD is a parallelogram. Find x and y using the following information: AB=4 (x+3) BC=6-(2+y) DC=12 (x-5) AD=3y

OpenStudy (anonymous):

Drawing a picture helps a lot  Since it's a parallelogram, you know opposite sides are going to be the same length, so: (1) AB = CD and (2) BC = AD For (1), plug in the values of AB and CD, so AB = CD ==> 4 (x+3) = 12 (x-5) solve for x, giving you ==> x = 9 For (2), plug in the values of BC and AD BC = AD ==> 6 - (2+y) = 3y solve for y, giving you ==> y = 1